# Should I Share this Translation? Evaluating Quality Feedback for User Reliance on Machine Translation

### v1

@@
As people increasingly use AI systems in work and daily life, feedback mechanisms that help them use AI responsibly are urgently needed, particularly in settings where users are not equipped to assess the quality of AI predictions. We study a realistic Machine Translation (MT) scenario where monolingual users decide whether to share an MT output, first without and then with quality feedback. We compare four types of quality feedback: explicit feedback that directly give users an assessment of translation quality using 1) error highlights and 2) LLM explanations, and implicit feedback that helps users compare MT inputs and outputs through 3) backtranslation and 4) question–answer (QA) tables. We find that all feedback types, except error highlights, significantly improve both decision accuracy and appropriate reliance. Notably, implicit feedback, especially QA tables, yields significantly greater gains than explicit feedback in terms of decision accuracy, appropriate reliance, and user perceptions – receiving the highest ratings for helpfulness and trust, and the lowest for mental burden. [^1] We release our code at https://github.com/dayeonki/mt_quality_feedback.
[…]
leftmargin=0pt, rightmargin=0pt, backgroundcolor=bggray, middlelinecolor=black, roundcorner=3
@@ 1 Introduction
Рис. 1: In our human study, each English-speaking monolingual participant reviews a sequence of 20 decision-making examples. Each example is shown in a two-step process: ➊ Independent decision-making: Participants first make judgments based solely on the English source and its Spanish MT output and ➋ AI-Assisted decision-making: They then reassess the same example with one of four randomly assigned feedback types . For each step, they respond to two questions: (1) Shareability: To the best of your knowledge, is the Spanish translation good enough to safely share with your Spanish-speaking neighbor? and (2) Confidence: How confident are you in your assessment?
@@
Prior studies have proposed quality feedback mechanisms to support MT decision-making, such as paraphrases, Quality Estimation (QE) scores, and backtranslation, but findings on their impact on user confidence and decision accuracy remain mixed (Zouhar et al., 2021; Mehandru et al., 2023). Building on this, we provide a more comprehensive assessment of which feedback types best support users in forming functional explanations to make reliable MT decisions. We evaluate four types of quality feedback, grouped by their mode of explicitness: 1) Explicit quality assessments of MT output (error highlights and LLM explanation); and 2) Implicit assessments that support input/output comparison (backtranslation and QA table).
@@ 3.1 Study Design
We study how different types of quality feedback impact users’ decision-making regarding MT shareability through a sequence of 20 examples in a between-subjects design, as illustrated in Figure 1. We present a scenario in which an English monolingual speaker reads official COVID-19 guidelines in English and decide whether the Spanish MT output is of sufficient quality to safely share with a Spanish-speaking neighbor. Each example is presented in two steps: ➊ Independent step, where participants first make judgments based solely on the English source and its Spanish translation, and a subsequent ➋ AI-Assisted step, where the same example is paired with a specific type of quality feedback to support decision-making. Examples were presented in randomized order, and two attention check questions were included.
@@
For each example, participants are first asked to make judgments based solely on an English sentence and its corresponding Spanish translation, without any quality feedback. They are asked to answer two questions: 1) Shareability: To the best of your knowledge, is the Spanish translation good enough to safely share with your Spanish-speaking neighbor? (with binary options: [Uncaptioned image] Safe to share as-is, and [Uncaptioned image] Needs bilingual review before sharing) [^2] We adopt the notion of shareability to capture perceived MT quality and the risk of incorrect communication, highlighting the potential consequences in high-stakes contexts.; and 2) Confidence: How confident are you in your assessment? (on a five-point Likert scale from 1:Very Unconfident to 5:Very Confident). Since the recruited participants are monolingual English speakers, they are instructed to do their best in assessing the shareability of the MT outputs, despite not being fluent in Spanish.
@@ 3.2 Types of Quality Feedback
Figure 2 illustrates an example of each quality feedback. While error highlights and LLM explanation provide explicit quality assessments of MT output, backtranslation and QA table provide implicit assessments that guide participants to compare MT input and output. Details on how each feedback is shown to participants are in Appendix C.1.333We compare the computational and time efficiency for generating each feedback in Appendix D.
@@
We generate natural language explanations using LLaMA-3 70b (Grattafiori et al., 2024). Instead of instructing the model to make a shareability decision, we prompt it to assess the overall quality of the Spanish MT relative to the English source text. The exact prompt is provided in Appendix A.1. For digestibility, we constrain the model to generate responses of fewer than three sentences. The generated explanations have an average length of 46.35 words.
@@
We use the AskQE framework (Ki et al., 2025) for question generation and answering, where questions are generated from the source text and answers are drawn from both the source and the backtranslated MT output. Specifically, we adopt an optimized version of AskQE that uses LLaMA-3 70b and entailed facts to guide question generation, and Google Translate for backtranslation of the Spanish MT output. All prompts are provided in Appendix A.2. On average, each example yields 2.65 questions, with an average question length of 10.04 words. As illustrated in Figure 2, we present the QA pairs in a table format with the following statements: ‘‘The questions are about the original English content’’ and ‘‘The answers are based on two sources: the original English text and the Spanish translation, which has been translated back into English for display’’. When the two answers are identical or highly similar, they are displayed in orange, else, blue. Similarity is computed using a soft variant of exact matching, with normalization for punctuation, case, whitespace, and articles.
@@ 3.3 Stimuli Collection
We sample 40 English-Spanish examples from the ContraTICO dataset (Ki et al., 2025), [^6] Criteria for data selection are detailed in Appendix B.2. which contains contrastive, synthetic MT errors in the COVID-19 domain. Each reference translation from the TICO-19 dataset (Anastasopoulos et al., 2020) is perturbed using eight linguistic perturbations, categorized as either minor or critical based on the potential real-world impact of the MT error. We recruit five bilingual annotators to independently annotate each MT output for gold shareability labels. Inter-annotator agreement, measured by Fleiss’ Kappa [^7] https://en.wikipedia.org/wiki/Fleiss_kappa, is moderate (0.449). We select 20 examples for the study based on high agreement scores determined by majority vote, with 10 examples per label. Further details are in Appendix B.1.
Рис. 3: Average decision accuracy (left) and CWA (right) for each type of quality feedback. *: significant with $p$-value < 0.05; **: $p$ < 0.01; ***: $p$ < 0.001; Non-marked ones are not statistically significant. Detailed results are provided in Appendix E.1.
@@ 3.4 Participants
We recruited 73 participants residing in the United States who self-identified English as their first, primary, and fluent language. Recruitment was conducted in two stages to exclude participants proficient in Spanish: 1) A pre-screening survey, where participants reported their English and Spanish proficiency on a five-point scale; and 2) The main task, limited to those who reported high English and low Spanish proficiency in the pre-screening phase. Each main task participant received 5 USD for completing the task (equivalent to 20 USD/hour), [^8] The task took a median of 15 minutes to complete. and 25 participants who achieved over 70% overall decision accuracy received an additional 2 USD bonus. Our institution’s IRB approved to conduct the study. Participants provided informed consent prior to the study. Further details are in Appendix C.2.
@@ 3.5 Dependent Variables
$\mathrm{sign}(\hat{s})=\begin{cases}1,&\text{if }\hat{s}=s^{*}\\ -1,&\text{otherwise}\end{cases}$
@@
Switch percentage is a widely used behavioral measure of reliance, capturing how often participants change their decisions after viewing AI feedback (Srivastava et al., 2022; He et al., 2023b). In our context, it reflects how quality feedback influences final shareability judgments (Eckhardt et al., 2024). We compute three metrics following the framework of Schemmer et al. (2023): 1) Over-reliance: the proportion of cases where a participant changes from a correct to an incorrect decision after feedback; 2) Under-reliance: the proportion of cases where a participant does not change from an incorrect decision to a correct one after the quality feedback; 3) Appropriate reliance: the proportion of cases where a participant either corrects an incorrect decision after receiving feedback (switch) or maintains a correct decision (no switch). To account for the differing consequences of reliance depending on shareability, we further break down each reliance metric by shareability label, as detailed in Appendix E.2.
@@ 4.1 RQ1: Does Quality Feedback Improve MT Decision-Making?
Рис. 4: Average decision accuracy (left) and CWA (right) for each type of quality feedback and shareability label. $\mathbf{n}$ indicates the number of examples aggregated for each condition and label. Independent aggregates responses made without quality feedback across all conditions. **: statistically significant with $p$-value < 0.01; ***: $p$ < 0.001; Non-marked ones are not statistically significant. Detailed results are provided in Appendix E.3.
@@ 4.2 RQ2: Which Feedback is Most Effective?
We perform pairwise independent Welch’s $t$-tests to compare different treatment conditions. Implicit feedback types generally outperform explicit ones. Participants who received QA table feedback have significantly higher AI-assisted decision accuracy (M = 0.765, S.E. = 0.022) than those in the error highlights (M = 0.647, S.E. = 0.025; p < 0.001) and the LLM explanation group (M = 0.693, S.E. = 0.024; p < 0.05), as shown in Figure 3. No significant difference was found between QA table and backtranslation (M = 0.716, S.E. = 0.025).
A similar pattern emerges for appropriate reliance (CWA). QA table group achieved significantly higher CWA (M = 0.488, S.E. = 0.040) than error highlights (M = 0.239, S.E. = 0.042; p < 0.001) and LLM explanations (M = 0.353, S.E. = 0.042; p < 0.05). Both backtranslation (M = 0.388, S.E. = 0.043; p < 0.05) and LLM explanation (p < 0.05) also yielded significantly higher CWA than error highlights. These results suggest that QA table feedback is the most effective overall, outperforming both explicit feedback types in supporting accurate and well-calibrated MT decisions.
@@ 4.3 RQ3: Which Feedback do Users Rely on Most?
We further examine switch percentages by shareability label in Appendix E.2. Participants show higher over-reliance and lower under-reliance for shareable examples than non-shareable ones. [^10] We refer to shareable as examples labeled “Safe to share as-is,” and non-shareable as those labeled “Needs bilingual review before sharing”. This suggests that participants are more likely to change their decisions when initially judging translations as shareable, but tend to maintain their decisions when judging them as non-shareable.
@@ 5 Analysis
5.3 What Makes Quality Feedback Helpful?
In contrast, participants who received implicit feedback described a more self-directed decision-making process. Backtranslation was considered helpful for verifying ‘‘some tiny unsure details’’ or checking whether ‘‘the core meaning of the original English text was preserved in the Spanish MT’’ by comparing the two English texts. Similarly, QA table encouraged participants to revisit the MT output when a mismatch was detected (‘‘if the statement is blue, I double check the phrase again’’), ‘‘compare the words and make determinations’’ themselves. Detailed comments are provided in Appendix E.4.
@@ 7 Limitation
Our human study focuses on monolingual source speakers who self-identified as proficient in English but not in Spanish, simulating a scenario in which non-English speakers encounter MT outputs of COVID-19 articles originally written in English. An important complementary study remains – evaluating with the monolingual target speakers. This would require modifying our current quality feedback setup: 1) Error highlights would continue to be displayed on the MT output; 2) LLM explanations would need to be presented in the target language; 3) Backtranslation would no longer be a suitable feedback; and 4) For QA table, questions would need to be generated from the backtranslated source, with answers derived from both the backtranslation and the (possibly perturbed) MT output. Future work can explore this variant through a human-centered study to assess how different forms of quality feedback influence MT decision-making for monolingual target speakers.
@@
The scope of our study is limited to the current experimental setup. While we focus on the English-Spanish language pair, motivated by Spanish being the most widely spoken non-English language in the United States (U.S. Census Bureau, 2022), our findings may not generalize to other language pairs. For consistency and fair comparison, we use the same model (LLaMA-3 70B) to generate both LLM explanations and QA table feedback, and the same MT system (Google Translate) to produce backtranslations and the backtranslated MT outputs used in QA generation. Our evaluation is also limited to four types of quality feedback: error highlights, LLM explanations, backtranslation, and QA table. This leaves open important questions around the effectiveness of other feedback types and combinations, which we leave to future work.

### v2

@@
As people increasingly use AI systems in work and daily life, feedback mechanisms that help them use AI responsibly are urgently needed, particularly in settings where users are not equipped to assess the quality of AI predictions. We study a realistic Machine Translation (MT) scenario where monolingual users decide whether to share an MT output, first without and then with quality feedback. We compare four types of quality feedback: explicit feedback that directly give users an assessment of translation quality using (1) error highlights and (2) LLM explanations, and implicit feedback that helps users compare MT inputs and outputs through (3) backtranslation and (4) question–answer (QA) tables. We find that all feedback types, except error highlights, significantly improve both decision accuracy and appropriate reliance. Notably, implicit feedback, especially QA tables, yields significantly greater gains than explicit feedback in terms of decision accuracy, appropriate reliance, and user perceptions – receiving the highest ratings for helpfulness and trust, and the lowest for mental burden. [^1] https://github.com/dayeonki/mt_quality_feedback
@@ 1 Introduction
Рис. 1: Overview of our study setup. In our human study, each English-speaking monolingual participant reviews a sequence of 20 decision-making examples. Each example is shown in a two-step process: ➊ Independent decision-making: Participants first make judgments based solely on the English source and its Spanish MT output and ➋ AI-Assisted decision-making: They then reassess the same example with one of four randomly assigned feedback types . For each step, they respond to two questions: (1) Shareability: To the best of your knowledge, is the Spanish translation good enough to safely share with your Spanish-speaking neighbor? and (2) Confidence: How confident are you in your assessment?
@@
Prior studies have proposed quality feedback mechanisms to support MT decision-making, such as paraphrases, Quality Estimation (QE) scores, and backtranslation, but findings on their impact on user confidence and decision accuracy remain mixed (Zouhar et al., 2021; Mehandru et al., 2023). Building on this, we provide a more comprehensive assessment of which feedback types best support users in forming functional explanations to make reliable MT decisions. We evaluate four types of quality feedback, grouped by their mode of explicitness: (1) Explicit quality assessments of MT output (error highlights and LLM explanation); and (2) Implicit assessments that support input/output comparison (backtranslation and QA table).
@@ 3.1 Study Design
We study how different types of quality feedback impact users’ decision-making regarding MT shareability through a sequence of 20 examples in a between-subjects design, as illustrated in Figure 1. We use the notion of shareability to capture not only perceived MT quality but also the potential risk of miscommunication, highlighting the potential consequences in high-stakes contexts. This framing aligns with how people make such decisions in practice: while the choice to share or not is often made implicitly in real-world use cases, our study makes this decision more explicit, yet still allows participants to make their own judgment as they naturally would.
Specifically, we situate participants in a scenario where an English monolingual speaker reads official COVID-19 guidelines in English and decide whether the Spanish MT output is of sufficient quality to safely share with a Spanish-speaking neighbor. Each example is presented in two steps: ➊ Independent step, where participants first make judgments based solely on the English source and its Spanish translation, and a subsequent ➋ AI-Assisted step, where the same example is paired with a specific type of quality feedback to support decision-making. Examples are shown in randomized order, and two attention checks were included.
@@
For each example, participants are first asked to make judgments based solely on an English sentence and its corresponding Spanish translation, without any quality feedback. They are asked to answer two questions: (1) Shareability: To the best of your knowledge, is the Spanish translation good enough to safely share with your Spanish-speaking neighbor? (with binary options: [Uncaptioned image] Safe to share as-is, and [Uncaptioned image] Needs bilingual review before sharing); and (2) Confidence: How confident are you in your assessment? (on a five-point Likert scale from 1:Very Unconfident to 5:Very Confident). Since the recruited participants are monolingual English speakers, they are instructed to do their best in assessing the shareability of the MT outputs, despite not being fluent in Spanish.
@@ 3.2 Types of Quality Feedback
Figure 2 illustrates an example of each quality feedback. Error highlights and LLM explanation provide explicit quality assessments of MT output, whereas backtranslation and QA table offer implicit assessments that guide participants to compare MT input and output. Details on how each feedback is shown to participants are in Appendix C.1. To control for feedback quality, we balance the error rates of feedback predictions across all types (Appendix D). [^2] We compare the computational and time efficiency for generating each feedback in Appendix E.
@@
We generate natural language explanations using LLaMA-3.3 70B (Grattafiori et al., 2024). [^4] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct Instead of instructing the model to make a shareability decision, we prompt it to assess the overall quality of the Spanish MT relative to the English source text. The exact prompt is provided in Appendix A.1. For digestibility, we constrain the model to generate responses of fewer than three sentences. The generated explanations have an average length of 46.35 words.
@@
We use the AskQE framework (Ki et al., 2025) for question generation and answering, where questions are generated from the source text and answers are drawn from both the source and the backtranslated MT output. Specifically, we adopt an optimized version of AskQE that uses LLaMA-3.3 70B and entailed facts to guide question generation, and Google Translate for backtranslation of the Spanish MT output. All prompts are provided in Appendix A.2. On average, each example yields 2.65 questions, with an average question length of 10.04 words. As illustrated in Figure 2, we present the QA pairs in a table format with the following statements: ‘‘The questions are about the original English content’’ and ‘‘The answers are based on two sources: the original English text and the Spanish translation, which has been translated back into English for display’’. When the two answers are identical or highly similar, they are displayed in orange, else, blue. Similarity is computed using a soft variant of exact matching, with normalization for punctuation, case, whitespace, and articles.
@@ 3.3 Stimuli Collection
We sample 40 English-Spanish examples from the ContraTICO dataset (Ki et al., 2025), [^6] Criteria for data selection are detailed in Appendix B.2. which contains contrastive, synthetic MT errors in the COVID-19 domain. Each reference translation from the TICO-19 dataset (Anastasopoulos et al., 2020) is perturbed using eight linguistic perturbations, categorized as either minor or critical based on the potential real-world impact of the MT error. We recruit five bilingual annotators to independently annotate each MT output for gold shareability labels. Inter-annotator agreement, measured by Fleiss’ Kappa [^7] https://en.wikipedia.org/wiki/Fleiss_kappa, is moderate (0.449). We select 20 examples for the study based on high agreement scores determined by majority vote, with 10 examples per label. Further details are in Appendix B.1.888English source sentences contain 471 words in total (average 23.6 words per sentence), and the Spanish translations contain 590 words in total (average 29.5 words per sentence).
Рис. 3: Average decision accuracy (left) and CWA (right) for each type of quality feedback. Paired-sample $t$-tests are performed to compare independent and AI-assisted performance and linear mixed-effects ANOVA with Bonferroni corrections to compare different treatment conditions. *: significant with $p$-value < 0.05; **: $p$ < 0.01; ***: $p$ < 0.001; Non-marked ones are not statistically significant. Detailed results are provided in Appendix G.1.
@@ 3.4 Participants
We recruited 73 participants residing in the United States who self-identified English as their first, primary, and fluent language. Recruitment was conducted in two stages to exclude participants proficient in Spanish: (1) A pre-screening survey, where participants reported their English and Spanish proficiency on a five-point scale; and (2) The main task, limited to those who reported high English and low Spanish proficiency in the pre-screening phase. Each main task participant received 5 USD for completing the task (equivalent to 20 USD/hour), [^9] The task took a median of 15 minutes to complete. and 25 participants who achieved over 70% overall decision accuracy received an additional 2 USD bonus. Our institution’s IRB approved to conduct the study. Participants provided informed consent prior to the study. Further details are in Appendix C.2.
@@ 3.5 Dependent Variables
[…]
@@
Switch percentage is a widely used behavioral measure of reliance, capturing how often participants change their decisions after viewing AI feedback (Srivastava et al., 2022; He et al., 2023b). In our context, it reflects how quality feedback influences final shareability judgments (Eckhardt et al., 2024). We compute three metrics following the framework of Schemmer et al. (2023): (1) Over-reliance: the proportion of cases where a participant changes from a correct to an incorrect decision after feedback; (2) Under-reliance: the proportion of cases where a participant does not change from an incorrect decision to a correct one after the quality feedback; (3) Appropriate reliance: the proportion of cases where a participant either corrects an incorrect decision after receiving feedback (switch) or maintains a correct decision (no switch). To account for the differing consequences of reliance depending on shareability, we further break down each reliance metric by shareability label, as detailed in Appendix G.2.
@@ 4 Results
Рис. 4: Average decision accuracy (left) and CWA (right) for each type of quality feedback and shareability label. $\mathbf{n}$ indicates the number of examples aggregated for each condition and label. Independent aggregates responses made without quality feedback across all conditions. **: statistically significant with $p$-value < 0.01; ***: $p$ < 0.001; Non-marked ones are not statistically significant. Detailed results are provided in Appendix G.3.
@@ 4.1 RQ1: Does Quality Feedback Improve MT Decision-Making?
In Appendix F, we further show that our two-step setup (➊ Independent → ➋ AI-assisted) does not induce learning effects. The observed gains in decision accuracy and appropriate reliance are attributable to the quality feedback, not repetition.
@@ 4.2 RQ2: Which Feedback is Most Effective?
We perform linear mixed-effects ANOVA, followed by Bonferroni correction for multiple comparisons across treatment conditions. Implicit feedback types generally outperform explicit ones. Participants who received QA table feedback have significantly higher AI-assisted decision accuracy (M = 0.765, S.E. = 0.022) than those in the error highlights (M = 0.647, S.E. = 0.043; p < 0.01) and the LLM explanation group (M = 0.693, S.E. = 0.042; p < 0.05), as shown in Figure 3. No significant difference was found between QA table and backtranslation (M = 0.716, S.E. = 0.025).
A similar pattern emerges for appropriate reliance (CWA). QA table group achieved significantly higher CWA (M = 0.488, S.E. = 0.040) than error highlights (M = 0.239, S.E. = 0.071; p < 0.01) and LLM explanations (M = 0.353, S.E. = 0.070; p < 0.05). Backtranslation (M = 0.388, S.E. = 0.072; p < 0.05) also yielded significantly higher CWA than error highlights. These results suggest that QA table feedback is the most effective overall, outperforming both explicit feedback types in supporting accurate and well-calibrated MT decisions.
@@ 4.3 RQ3: Which Feedback do Users Rely on Most?
We further examine switch percentages by shareability label in Appendix G.2. Participants show higher over-reliance and lower under-reliance for shareable examples than non-shareable ones. [^11] We refer to shareable as examples labeled “Safe to share as-is,” and non-shareable as those labeled “Needs bilingual review before sharing”. This suggests that participants are more likely to change their decisions when initially judging translations as shareable, but tend to maintain their decisions when judging them as non-shareable.
@@ 5 Analysis
5.3 What Makes Quality Feedback Helpful?
In contrast, participants who received implicit feedback described a more self-directed decision-making process. Backtranslation was considered helpful for verifying ‘‘some tiny unsure details’’ or checking whether ‘‘the core meaning of the original English text was preserved in the Spanish MT’’ by comparing the two English texts. Similarly, QA table encouraged participants to revisit the MT output when a mismatch was detected (‘‘if the statement is blue, I double check the phrase again’’), ‘‘compare the words and make determinations’’ themselves. Detailed comments are provided in Appendix G.4.
@@ 7 Limitation
Presentation differences.
We acknowledge that presentation differences, such as formatting or visual salience, can influence user behavior independently of the underlying feedback type. To mitigate this, we iteratively designed and piloted each feedback interface to ensure clarity and minimize usability discrepancies across conditions. Specifically, we conducted seven rounds of internal pilot testing and usability checks to identify sources of confusion, refine language, or layout. While some format-specific differences were necessary (e.g., error highlights inherently rely on color to convey span-level quality signals, while QA tables require side-by-side comparisons), we carefully calibrated to ensure that each condition represented a ‘‘best of its breed’’ version of that feedback type. Therefore, although we acknowledge the possibility of minor usability-related effects, we believe these differences are unlikely to fully account for the performance gaps observed between feedback types.
Study with monolingual target speakers.
@@
Our human study focuses on monolingual source speakers who self-identified as proficient in English but not in Spanish, simulating a scenario in which non-English speakers encounter MT outputs of COVID-19 articles originally written in English. An important complementary study remains – evaluating with the monolingual target speakers. This would require modifying our current quality feedback setup: (1) Error highlights would continue to be displayed on the MT output; (2) LLM explanations would need to be presented in the target language; (3) Backtranslation would no longer be a suitable feedback; and (4) For QA table, questions would need to be generated from the backtranslated source, with answers derived from both the backtranslation and the (possibly perturbed) MT output. Future work can explore this variant through a human-centered study to assess how different forms of quality feedback influence MT decision-making for monolingual target speakers.
Limited scope.
@@
The scope of our study is limited to the current experimental setup. While we focus on the English-Spanish language pair, motivated by Spanish being the most widely spoken non-English language in the United States (U.S. Census Bureau, 2022), our findings may not generalize to other language pairs. For consistency and fair comparison, we use the same model (LLaMA-3.3 70B) to generate both LLM explanations and QA table feedback, and the same MT system (Google Translate) to produce backtranslations and the backtranslated MT outputs used in QA generation. Our evaluation is also limited to four types of quality feedback: error highlights, LLM explanations, backtranslation, and QA table.
Moreover, we intentionally focused on a single domain (communicating COVID-19 protocols) and a specific user population with specific language background in order to ensure that our study reflects realistic decision-making scenarios. This controlled setting helped participants better operationalize the notion of shareability and limited potential confounding factors in the study. We believe the core findings ought to generalize to other language pairs since MT quality was not a major factor, and possibly other domains with similar notions of shareability.
Understanding the effect of quality feedback on MT reliance across a wider range of contexts is an important direction of future work. For instance, the doctor-patient setting represents a valuable use case where the doctor as a user of MT has more consideration and technical expertise in deciding whether MT outputs are shareable. Another example could be dialogues, where additional rounds of human and automatic feedback is possible. We view our study as a foundational step toward more extensive future research in this area.
Use of synthetic dataset.
While our study aimed to simulate a realistic decision-making scenario, the MT outputs themselves are drawn from a synthetically constructed dataset ContraTICO. This design choice allowed us to systematically control for specific error types and severity levels. However, we acknowledge that these errors may not fully capture the complexity and variability of naturally occurring MT errors.